OpenStudy (anonymous):

integrate sqrt(1-cosx)

OpenStudy (anonymous):

\[-2\sqrt{1-\cos (x )} \cot(x/2) + constant\]

OpenStudy (anonymous):

i got 2sqrt2 cosx/2 ans. is this correct or not?

OpenStudy (anonymous):

Its wrong.

OpenStudy (anonymous):

isnt it \[-2\sqrt{1+\cos x}+c\]

OpenStudy (anonymous):

No that's wrong, you forgot to put a cotangent function of 0.5x in there and also its negative cosine of x.

OpenStudy (anonymous):

\[\sqrt{1-(1-2\sin^2x/2)}\][]\sqrt{2}\] i did this.is this wrong??